# On the Hausdorff dimension and Cantor set structure of sliding Shilnikov invariant sets

### v2

@@ 1.3. Main result
Let $V\subseteq\mathbb{R}^{3}$ be an open subset and consider a Filippov system $Z=(X,Y)\in\Omega_{g}^{1,1}(V,\mathbb{R}^{3})$ possessing a sliding Shilnikov connection $\Gamma$ passing through a visible fold-regular point $q\in\partial M^{s,e}$. Consider the first return map $\pi:\operatorname{Dom}(\pi)\subseteq\gamma\to\gamma$ associated to $\Gamma$. Then, there exists a neighbourhood $U\subseteq\gamma$ of $q$ such that:
@@
Recall that a metric space $K\neq\emptyset$ is called a Cantor set if it is a compact, totally disconnected (the only connected subsets are singletons), and perfect (every point is an accumulation point). It is well-known that any non-empty metric spaces with these topological properties are homeomorphic to each other (see, for example, [25, Chapter 2, Theorem 67 and Theorem 73]).
Theorem A is proven in Section 3. Its proof is based on iterated function systems theory, which is introduced in Section 2.
@@ 2.1. Hausdorff Dimension
The Hausdorff dimension is a notion of dimension related to the “size”, in the measure sense, of a set. More specifically, let us consider the diameter of an arbitrary subset $V\subseteq\mathbb{R}^{d}$,
@@
Given a set $S$ and a real number $s\geq 0$, for each $\delta>0$ we look for all countable (possibly finite) covers of $S$ with subsets of diameter at most $\delta$ (also called a $\delta$-cover) and calculate the quantity
@@
We define the $s$-dimensional Hausdorff measure of $S$ as the limit $\cal{H}^{s}(S):=\lim_{\delta\to 0}\cal{H}^{s}_{\delta}(S)$. The Hausdorff dimension of $S$, denoted by $\operatorname{dim_{H}}\left(S\right)$, is then determined by
@@
    If $S_{i}\subseteq S_{j}$, then $\operatorname{dim_{H}}\left(S_{i}\right)\leq\operatorname{dim_{H}}\left(S_{j}\right)$ ([27, Theorem 2(1)]);
@@
    If $\left\{S_{k}\right\}_{k\geq 1}$ is a countable collection of sets, then $\operatorname{dim_{H}}\left(\bigcup_{k\geq 1}S_{k}\right)=\sup_{k\geq 1}\left\{\operatorname{dim_{H}}\left(S_{k}\right)\right\}$ ([27, Theorem 2(2)]);
@@
    If $f$ is a bi-Lipschitz map, then $\operatorname{dim_{H}}\left(S\right)=\operatorname{dim_{H}}\left(f(S)\right)$, which implies that diffeomorphic compact sets have the same Hausdorff dimension ([27, Theorem 2(5)]);
@@
    If $\operatorname{dim_{H}}\left(S\right)<1$, then its one-dimensional Lebesgue measure is 0 ([27, Theorem 2(8)]);
@@ 2.2. Iterated Function Systems (IFS)
A key concept concerning an IFS is its attractor set. To define it, let $p:I^{\mathbb{N}}\to K$ be the projection map from the symbol space $I^{\mathbb{N}}$ into $K$, that is, for each $\omega=(\omega_{1},\omega_{2},\ldots)\in I^{\mathbb{N}}$, $p(\omega)$ is defined satisfying
$\left\{p(\omega)\right\}=\bigcap_{k\geq 1}f_{\omega|_{k}}(K),$ (4)
where $\omega|_{k}:=(\omega_{1},\ldots,\omega_{k})$, for $k\geq 1$. Then, the attractor set of the IFS is defined as
$\Delta:=p\left(I^{\mathbb{N}}\right)=\bigcup_{\omega\in I^{\mathbb{N}}}\bigcap_{k\geq 1}f_{\omega|_{k}}(K).$ (5)
@@ 2.3. Conformal Iterated Function Systems (CIFS)
Let :=$\mathcal{F}:=\left\{f_{i}:K\subseteq\mathbb{R}^{d}\to K\right\}_{i\in I}$ be an IFS. In what follows, we list the conformal conditions:
@@
    For each $i\in I$, the function $f_{i}$ is an injection of $K$ into itself;
@@
    $\left\lvert{f_{i}(x)-f_{i}(y)}\right\rvert\leq s\left\lvert{x-y}\right\rvert,\quad\text{for all }i\in I\text{ and for all }x,y\in K;$
@@
    (Open Set Condition) The set $K$ is connected, and each function satisfies $f_{i}(\operatorname{Int}(K))\subseteq\operatorname{Int}(K)$ and $f_{i}(\operatorname{Int}(K))\cap f_{j}(\operatorname{Int}(K))=\emptyset$, for all $i,j\in I,i\neq j$;
@@
    There is an open set $V\subseteq\mathbb{R}^{d}$, with $K\subseteq V$, such that each $f_{i}$ extends to $f_{i}^{V}$, a $\cal{C}^{1,\varepsilon}$ diffeomorphism on $V$, and the extensions are conformal functions, that is, the derivatives $Df_{i}^{V}(x)$ satisfy $Df_{i}^{V}(x)=\kappa_{x,i}\operatorname{Isom}_{x,i}$, where $\kappa_{x,i}\in\mathbb{R}$ and $\operatorname{Isom}_{x,i}:\mathbb{R}^{d}\to\mathbb{R}^{d}$ is an isometry, for any $x\in V$ and $i\in I$ (for more details on conformal functions, see [23, Chapter 7]);
@@
    where $\operatorname{m}_{d}$ is the $d$-dimensional Lebesgue measure;
@@
    where $f_{i}^{V}$, $i\in I$, are the extensions establish in (C4), $x,y\in V$, $\left\lVert{\cdot}\right\rVert_{\operatorname{op}}$ is the operator norm, and $\left\lVert{(Df_{i}^{V})^{-1}}\right\rVert_{\operatorname{unif}}:=\sup_{z\in V}\left\{\left\lVert{(Df_{i}^{V}(z))^{-1}}\right\rVert_{\operatorname{op}}\right\}$.
An IFS that satisfies all these conditions is said to be a conformal IFS or just a CIFS.
@@
Let =$\mathcal{F}=\left\{f_{i}\right\}_{i\in I}$ be a CIFS with attractor set $\Delta$. Then, the identity
@@
holds, where $\operatorname{Fin}(I)\subseteq 2^{I}$ is the set of finite subsets of $I$.
@@
Now we present another important result concerning CIFS, which states that in the definition of the attractor set (5), the union and intersection can be interchanged.
@@ 3. Proof of the main result
The local dynamics of a sliding Shilnikov connection is captured by the first-return map $\pi$, defined on a subset of $\partial M^{s,e}$ within a small neighborhood $U\subseteq\partial M^{s,e}$ of the visible-fold regular point $q$. Thus, our initial task is to construct this map. Subsequently, based on its construction, we will relate this map to the theory of CIFS by showing that it can be branched into countably many smooth maps such that the functions corresponding to their inverses constitute a CIFS.
@@ 3.1. Construction of the First Return Map
Let us consider a Shilnikov connection $\Gamma$ (see Definition A) with a hyperbolic pseudo saddle-focus $p\in M^{s}$ and passing through a visible fold-regular point $q\in\partial M^{s}$, which is a visible fold point of $X$. Recall that, for $u\in V\subseteq\mathbb{R}^{3}$ and $w\in\overline{M^{s}}$, the flows of $X$ and $\tilde{Z}$ are denoted, respectively, by $\varphi^{t}_{X}(u)$ and $\varphi^{t}_{\tilde{Z}}(w)$.
@@
Property (A2) from Remark 1 implies that, for sufficiently small $r>0$, $\gamma_{r}$ is a smooth curve of visible fold-regular points of $Z$, which are also visible fold points for $X$. In addition, Property (A1) from Remark 1 states that $\gamma_{r}$ is a transversal section of the sliding vector field $\tilde{Z}$.
@@
implying that $\mu_{r}$ is a smooth curve containing $p$. Furthermore, since $p\in\mu_{r}\subseteq M^{s}$ is a hyperbolic focus of $\tilde{Z}$, we can see $\mu_{r}\setminus{p}$ is transversal to $\tilde{Z}$. Indeed, if this were not the case, the vector $0\neq v\in T_{p}M$ tangent to $\mu_{r}$ at $p$ would be a real eigenvector of $D\tilde{Z}(p)$ restricted to $T_{p}M$. However, this is impossible because the restriction of $D\tilde{Z}(p)$ to $T_{p}M$ has a pair of complex conjugate eigenvalues.
@@
From Definition A, we have $\lim_{t\to-\infty}\varphi^{t}_{\tilde{Z}}(q)=p$. Therefore, taking into account that $\gamma_{r}$ is a transversal section of the $\tilde{Z}$ and that $p\in\mu_{r}\subseteq M^{s}$ is a hyperbolic focus of $\tilde{Z}$, the Implicit Function Theorem can be used to ensure that, for sufficiently small $r>0$, $\varphi^{t}_{\tilde{Z}}(w)$ converges to $p$ as $t\to-\infty$ for all $w\in\gamma_{r}$. Now, since $\mu_{r}\setminus\{p\}$ is transversal to $\tilde{Z}$ and $p\in\mu_{r}\subseteq M^{s}$ is a hyperbolic focus of $\tilde{Z}$, we deduce that
@@
where the collections $\left\{L_{i}^{\mu_{r}}\right\}_{i\geq 0},\left\{R_{i}^{\mu_{r}}\right\}_{i\geq 0}\subseteq\mu_{r}$, of the connect components of $S_{r}\cap\mu_{r}$, are located in opposite sides of $p\in\mu_{r}$, $L_{i}^{\mu_{r}}\cap L_{j}^{\mu_{r}}=\emptyset=R_{i}^{\mu_{r}}\cap R_{j}^{\mu_{r}}$, if $i\neq j$, and $L_{i}^{\mu_{r}},R_{i}^{\mu_{r}}$ converge to $\left\{p\right\}$ in Hausdorff distance as $i$ increases (see Figure 3). In addition, for each $\zeta\in S_{r}\cap\mu_{r}$, there exists $t_{\tilde{Z}}(\zeta)>0$ satisfying $\varphi^{t_{\tilde{Z}}(\zeta)}_{\tilde{Z}}(\zeta)\in\gamma_{r}$ and $\varphi^{t}_{\tilde{Z}}(\zeta)\in M^{s}$, for all $t\in(0,t_{\tilde{Z}}(\zeta))$. This induces the following maps from each $J\in\{L_{i}^{\mu_{r}}\}_{i\geq 0}\cup\{R_{i}^{\mu_{r}}\}_{i\geq 0}$ into $\gamma_{r}$,
@@
For $J\in\{L_{i}^{\mu_{r}}\}_{i\geq 1}\cup\{R_{i}^{\mu_{r}}\}_{i\geq 1}$, such maps are diffeomorphism. However, for $J\in\{L_{0}^{\mu_{r}},R_{0}^{\mu_{r}}\}$, the induced map $\theta_{\tilde{Z}}^{J}$ may not be surjective (see Figure 3).
@@
where $\theta_{X}$ is the diffeomorphism given by (6). This ensures that the properties $L_{i}^{\gamma_{r}}\cap L_{j}^{\gamma_{r}}=\emptyset$ and $R_{i}^{\gamma_{r}}\cap R_{j}^{\gamma_{r}}=\emptyset$ hold for $i\neq j$, and that $L_{i}^{\gamma_{r}},R_{i}^{\gamma_{r}}$ converge to the $\{q\}$ in Hausdorff distance as $i$ increases. Thus, we define
@@ 3.2. The First Return Map as a CIFS
For each $J\in\cal{J}_{U}$, the restriction $\pi_{J}=:\pi|_{J}:J\to\gamma_{r}$ is a diffeomorphims, thus its inverse $\psi_{J}:=\pi_{J}^{-1}:\gamma_{r}\to J$ is well-defined. In what follows we will investigate the behavior of $\pi_{J}$ and $\psi_{J}$.
@@
Now, in the definition of the first return map (8), take $r\in(0,R_{0})$ such that $\mu_{r}\subseteq\cal{O}$. Denote by $\rho_{X}$ the restriction of $\theta_{X}$, given by (6), to $\cal{W}_{r}$, which is a diffeomorphims onto its image $S_{r}\cap\mu_{r}$. In addition, given $w\in\cal{W}_{r}$, since $\rho_{X}(w)\in S_{r}\cap\mu_{r}\subseteq\cal{O}$, we can define an integer $c(w)$ corresponding to the number of times that $\rho$ must be applied to $\rho_{X}(w)$ before it enters $L_{1}^{\mu_{r}}\cup R_{1}^{\mu_{r}}$, that is,
@@
$\rho_{\tilde{Z}}^{J}:=\begin{cases}\theta_{\tilde{Z}}^{L_{1}^{\mu_{r}}}(\zeta)&\text{if }J\in\{L_{i}^{\gamma_{r}}\}_{i\geq 0},\vspace{0.15cm}\\ \theta_{\tilde{Z}}^{R_{1}^{\mu_{r}}}(\zeta)&\text{if }J\in\{R_{i}^{\gamma_{r}}\}_{i\geq 0},\\ \end{cases}$
Notice that, for $J\in\{L_{i}^{\gamma_{r}}\}_{i\geq 0}$, we have $\rho_{\tilde{Z}}^{J}|_{L_{1}^{\mu_{r}}}=\theta_{\tilde{Z}}^{L_{1}^{\mu_{r}}}$, and for $J\in\{R_{i}^{\gamma_{r}}\}_{i\geq 0}$, we have $\rho_{\tilde{Z}}^{J}|_{R_{1}^{\mu_{r}}}=\theta_{\tilde{Z}}^{R_{1}^{\mu_{r}}}$. These maps are diffeomorphisms, respectively, from $L_{1}^{\mu_{r}}$ and $R_{1}^{\mu_{r}}$ onto $\gamma_{r}$, induced by $\tilde{Z}$ as defined in (7). It is important to mention that the guaranteed diffeomorphic nature of these maps is the reason we chose $L_{1}^{\mu_{r}}\cup R_{1}^{\mu_{r}}$ to define $c(w)$, rather than $L_{0}^{\mu_{r}}\cup R_{0}^{\mu_{r}}$.
@@
There exists a neighborhood $U\subseteq\gamma_{r}$ of $q$ satisfying (P) for which $10<\epsilon_{J}\leq\left\lvert{\psi^{\prime}_{J}(x)}\right\rvert\leq s<$ 1, for every $J\in\cal{J}_{U}$ and for every $x\in\gamma_{r}$, and, consequently, $\left\lvert{\pi^{\prime}(w)}\right\rvert\geq 1/s>1$, for every $w\in U\cap\cal{W}_{r}$.
@@
Now, since $\rho_{X}$, $\rho_{\tilde{Z}}^{J}$, and $H|_{\mu_{r}}$ are diffeomorphisms on compact sets, we can set $\alpha_{X}:=\min_{x\in\gamma_{r}}\left\lvert{\rho_{X}^{\prime}(x)}\right\rvert>0$, $\alpha_{\tilde{Z}}:=\min_{\zeta\in J}\big{|}(\rho_{\tilde{Z}}^{J})^{\prime}(\zeta)\big{|}>0$, $\alpha_{H}:=\min_{\zeta\in\mu_{r}}\left\lvert{H(\zeta)}\right\rvert>0$, and $\alpha_{H^{-1}}:=\min_{z\in H(\mu_{r})}\left\lvert{H^{-1}(z)}\right\rvert>0$, and define $A:=\alpha_{X}\alpha_{\tilde{Z}}\alpha_{H}\alpha_{H^{-1}}>0$. Thus, taking the expression (10) into account, we get
@@
We need to clarify an important detail. Formally, the functions of the IFS $\Psi$ are defined on the curve $\gamma_{r}$, but working with them in this context can be cumbersome. Instead, we can consider an orientation-preserving diffeomorphism $h:\gamma_{r}\to[-1,1]$ such that $h(q)=0$. By defining $\tilde{\psi}_{J}:=h\circ\psi_{J}\circ h^{-1}:[-1,1]\to h(J)$, we form the set $\tilde{\Psi}:=\big{\{}\tilde{\psi}_{J}:J\in\cal{J},J\subseteq U\big{\}}$, which is an IFS. Additionally, we have $h(L_{i}^{\gamma_{r}})\subseteq[-1,0)$ and $h(R_{i}^{\gamma_{r}})\subseteq(0,1]$. This provides a nice graphical representation of the first return map $\pi$ in Figure 4. Since Property (B3) implies that the Hausdorff dimension is invariant under bi-Lipschitz maps, which is the case for $h$, this transformation does not change any of the properties we are investigating. Therefore, from now on, we will assume, by an abuse of notation, that the functions in $\Psi$ are defined in the interval $[-1,1]$ and that the sets $J\in\cal{J}$ are subintervals of $[-1,1]$.

### v4

@@ 1.3. Main result
Let $V\subseteq\mathbb{R}^{3}$ be an open subset and consider a Filippov system $Z=(X,Y)\in\Omega_{g}^{1,1}(V,\mathbb{R}^{3})$ possessing a sliding Shilnikov connection $\Gamma$ passing through a visible fold-regular point $q\in\partial M^{s,e}$. Consider the first return map $\pi:\operatorname{Dom}(\pi)\subseteq\gamma\to\gamma$ associated to $\Gamma$. Then, there exists a neighborhood $U\subseteq\gamma$ of $q$ such that:
@@
Recall that a metric space $K\neq\emptyset$ is called a Cantor set if it is compact, totally disconnected (the only connected subsets are singletons), and perfect (every point is an accumulation point). It is well-known that all non-empty metric spaces with these topological properties are homeomorphic to each other (see, for example, [25, Chapter 2, Theorem 67 and Theorem 73]).
The proof of Theorem A is made in Section 3, and is based on iterated function systems theory, which is introduced in Section 2.
@@ 2.1. Hausdorff Dimension
The Hausdorff dimension is a notion of dimension related to the “size”, in a measure sense, of a set. More precisely, let us consider the diameter of an arbitrary subset $V\subseteq\mathbb{R}^{d}$,
@@
Given a set $S$ and a real number $s\geq 0$, for each $\delta>0$ we look at all countable (possibly finite) covers of $S$ with subsets of diameter at most $\delta$ (also called a $\delta$-cover) and calculate the quantity
@@
We define the $s$-dimensional Hausdorff measure of $S$ as the limit
$\cal{H}^{s}(S):=\lim_{\delta\to 0}\cal{H}^{s}_{\delta}(S).$
The Hausdorff dimension of $S$, denoted by $\operatorname{dim_{H}}\left(S\right)$, is then determined by
@@
    If $S_{i}\subseteq S_{j}$, then $\operatorname{dim_{H}}\left(S_{i}\right)\leq\operatorname{dim_{H}}\left(S_{j}\right)$ ([27, Theorem 2(1)]).
@@
    If $\left\{S_{k}\right\}_{k\geq 1}$ is a countable collection of sets, then $\operatorname{dim_{H}}\left(\bigcup_{k\geq 1}S_{k}\right)=\sup_{k\geq 1}\left\{\operatorname{dim_{H}}\left(S_{k}\right)\right\}$ ([27, Theorem 2(2)]).
@@
    If $f$ is a bi-Lipschitz map, then $\operatorname{dim_{H}}\left(S\right)=\operatorname{dim_{H}}\left(f(S)\right)$, which implies that diffeomorphic compact sets have the same Hausdorff dimension ([27, Theorem 2(5)]).
@@
    If $\operatorname{dim_{H}}\left(S\right)<1$, then the one-dimensional Lebesgue measure of $S$ is 00 ([27, Theorem 2(8)]).
@@ 2.2. Iterated Function Systems (IFS)
A key concept concerning an IFS is its attractor set. To define it, let $\operatorname{proj}:I^{\mathbb{N}}\to 2^{K}$ be the projection map from the symbol space $I^{\mathbb{N}}$ into $K$, that is, for each $\omega=(\omega_{1},\omega_{2},\ldots)\in I^{\mathbb{N}}$, $\operatorname{proj}(\omega)$ is defined satisfying
$\operatorname{proj}(\omega)=\bigcap_{k\geq 1}f_{\omega|_{k}}(K),$ (5)
where $\omega|_{k}:=(\omega_{1},\ldots,\omega_{k})$, for $k\geq 1$.
Then, the attractor set of the IFS is defined as
$\Delta:=\operatorname{proj}\left(I^{\mathbb{N}}\right)=\bigcup_{\omega\in I^{\mathbb{N}}}\bigcap_{k\geq 1}f_{\omega|_{k}}(K).$ (6)
@@
Remark 2.
It is worth noting that if there exists $10<s<$ 1 for which the contraction constants satisfy $c_{i}\leq s$, the set in (5) becomes a singleton for each $\omega\in I^{\mathbb{N}}$. In this case, $\operatorname{proj}$ can also be regarded as a function from $I^{\mathbb{N}}$ into $K$, in a minor abuse of notation.
@@ 2.3. Conformal Iterated Function Systems (CIFS)
An IFS $\cal{F}$ is said to be a conformal IFS, or just a CIFS, if it satisfies the following conditions:
@@
    For each $i\in I$, the function $f_{i}$ is an injection of $K$ into itself.
@@
    $\left\lvert{f_{i}(x)-f_{i}(y)}\right\rvert\leq s\left\lvert{x-y}\right\rvert,\quad\text{for all }i\in I\text{ and for all }x,y\in K.$
@@
    (Open Set Condition) The set $K$ is connected, and each function satisfies $f_{i}(\operatorname{Int}(K))\subseteq\operatorname{Int}(K)$ and $f_{i}(\operatorname{Int}(K))\cap f_{j}(\operatorname{Int}(K))=\emptyset$, for all $i,j\in I,i\neq j$.
@@
    There is an open set $V\subseteq\mathbb{R}^{d}$, with $K\subseteq V$, such that each $f_{i}$ extends to $f_{i}^{V}$, a $\cal{C}^{1,\varepsilon}$ diffeomorphism on $V$, and the extensions are conformal functions, that is, the derivatives $Df_{i}^{V}(x)$ satisfy $Df_{i}^{V}(x)=\kappa_{x,i}\operatorname{Isom}_{x,i}$, where $\kappa_{x,i}\in\mathbb{R}$ and $\operatorname{Isom}_{x,i}:\mathbb{R}^{d}\to\mathbb{R}^{d}$ is an isometry, for any $x\in V$ and $i\in I$ (for more details on conformal functions, see [23, Chapter 7]).
@@
    where $\operatorname{m}_{d}$ is the $d$-dimensional Lebesgue measure.
@@
    where $f_{i}^{V}$, $i\in I$, are the extensions established in (C4), $x,y\in V$, $\left\lVert{\cdot}\right\rVert_{\operatorname{op}}$ is the operator norm, and $\left\lVert{(Df_{i}^{V})^{-1}}\right\rVert_{\operatorname{unif}}:=\sup_{z\in V}\left\{\left\lVert{(Df_{i}^{V}(z))^{-1}}\right\rVert_{\operatorname{op}}\right\}$.
@@
Let =$\mathcal{F}=\left\{f_{i}\right\}_{i\in I}$ be a CIFS with attractor set $\Delta$. Then, the equality
@@
holds, where $\operatorname{Fin}(I)\subseteq 2^{I}$ is the collection of finite subsets of $I$.
@@
Now we present another proposition concerning CIFS, which states that in the definition of the attractor set (6), the union and intersection can be, in a sense, interchanged.
@@ 3. Proof of the main result
The local dynamics of a sliding Shilnikov connection is captured by the first return map $\pi$, defined on a subset of $\partial M^{s,e}$ within a small neighborhood $U\subseteq\partial M^{s,e}$ of the visible fold-regular point $q$. Thus, our initial task is to construct this map. Subsequently, based on its construction, we will relate this map to the theory of CIFS by showing that it can be branched into countably many smooth maps such that the functions corresponding to their inverses constitute a CIFS.
@@ 3.1. Construction of the First Return Map
Let us consider a sliding Shilnikov connection $\Gamma$ (see Definition A) with a hyperbolic pseudo-saddle-focus $p\in M^{s}$ and passing through a visible fold-regular point $q\in\partial M^{s}$, which is a visible fold point of $X$. Recall that, for $u\in V\subseteq\mathbb{R}^{3}$ and $w\in\overline{M^{s}}$, the flows of $X$ and $\tilde{Z}$ are denoted, respectively, by $\varphi^{t}_{X}(u)$ and $\varphi^{t}_{\tilde{Z}}(w)$.
@@
(A2) from Remark 1 implies that, for sufficiently small $r>0$, $\gamma_{r}$ is a smooth curve of visible fold-regular points of $Z$, which are also visible fold points for $X$. In addition, (A1) from Remark 1 states that $\gamma_{r}$ is a transversal section of the sliding vector field $\tilde{Z}$.
@@
implying that $\mu_{r}$ is a smooth curve containing $p$. Furthermore, since $p\in\mu_{r}\subseteq M^{s}$ is a hyperbolic focus of $\tilde{Z}$, we can see that $\mu_{r}\setminus\left\{p\right\}$ is transversal to $\tilde{Z}$. Indeed, if this was not the case, the vector $0\neq v\in T_{p}M$ tangent to $\mu_{r}$ at $p$ would be a real eigenvector of $D\tilde{Z}(p)$ restricted to $T_{p}M$. However, this is impossible because the restriction of $D\tilde{Z}(p)$ to $T_{p}M$ has a pair of complex conjugate eigenvalues.
@@
From Definition A, we have $\lim_{t\to-\infty}\varphi^{t}_{\tilde{Z}}(q)=p$. Therefore, taking into account that $\gamma_{r}$ is a transversal section of $\tilde{Z}$ and that $p\in\mu_{r}\subseteq M^{s}$ is a hyperbolic focus of $\tilde{Z}$, the Implicit Function Theorem can be used to ensure that, for sufficiently small $r>0$, $\varphi^{t}_{\tilde{Z}}(w)$ converges to $p$ as $t\to-\infty$ for all $w\in\gamma_{r}$. Now, since $\mu_{r}\setminus\{p\}$ is transversal to $\tilde{Z}$ and $p\in\mu_{r}\subseteq M^{s}$ is a hyperbolic focus of $\tilde{Z}$, we deduce that
@@
where the collections $\left\{L_{i}^{\mu_{r}}\right\}_{i\geq 0},\left\{R_{i}^{\mu_{r}}\right\}_{i\geq 0}\subseteq\mu_{r}$, of the connected components of $S_{r}\cap\mu_{r}$, are located in opposite sides of $p\in\mu_{r}$, $L_{i}^{\mu_{r}}\cap L_{j}^{\mu_{r}}=\emptyset=R_{i}^{\mu_{r}}\cap R_{j}^{\mu_{r}}$, if $i\neq j$, and $L_{i}^{\mu_{r}},R_{i}^{\mu_{r}}$ converge to $\left\{p\right\}$ in the Hausdorff distance as $i$ increases (see Figure 3). In addition, for each $\zeta\in S_{r}\cap\mu_{r}$, there exists $t_{\tilde{Z}}(\zeta)>0$ satisfying $\varphi^{t_{\tilde{Z}}(\zeta)}_{\tilde{Z}}(\zeta)\in\gamma_{r}$ and $\varphi^{t}_{\tilde{Z}}(\zeta)\in M^{s}$, for all $t\in(0,t_{\tilde{Z}}(\zeta))$. This induces the following maps from each $J\in\{L_{i}^{\mu_{r}}\}_{i\geq 0}\cup\{R_{i}^{\mu_{r}}\}_{i\geq 0}$ into $\gamma_{r}$,
@@
For $J\in\{L_{i}^{\mu_{r}}\}_{i\geq 1}\cup\{R_{i}^{\mu_{r}}\}_{i\geq 1}$, such maps are diffeomorphisms. However, for $J\in\{L_{0}^{\mu_{r}},R_{0}^{\mu_{r}}\}$, the induced map $\theta_{\tilde{Z}}^{J}$ may not be surjective (see Figure 3).
@@
where $\theta_{X}$ is the diffeomorphism given by (7). This ensures that the properties $L_{i}^{\gamma_{r}}\cap L_{j}^{\gamma_{r}}=\emptyset$ and $R_{i}^{\gamma_{r}}\cap R_{j}^{\gamma_{r}}=\emptyset$ hold for $i\neq j$, and that $L_{i}^{\gamma_{r}},R_{i}^{\gamma_{r}}$ converge to $\{q\}$ in the Hausdorff distance as $i$ increases. Thus, we define
@@ 3.2. The First Return Map as a CIFS
For each $J\in\cal{J}_{U}$, the restriction $\pi_{J}:=\pi|_{J}:J\to\gamma_{r}$ is a diffeomorphism, thus its inverse $\psi_{J}:=\pi_{J}^{-1}:\gamma_{r}\to J$ is well-defined. In what follows we will investigate the behavior of $\pi_{J}$ and $\psi_{J}$.
@@
Now, in the definition of the first return map (9), take $r\in(0,R_{0})$ such that $\mu_{r}\subseteq\cal{O}$. Denote by $\rho_{X}$ the restriction of $\theta_{X}$, given by (7), to $\cal{W}_{r}$, which is a diffeomorphism onto its image $S_{r}\cap\mu_{r}$. In addition, given $w\in\cal{W}_{r}$, since $\rho_{X}(w)\in S_{r}\cap\mu_{r}\subseteq\cal{O}$, we can define an integer $c(w)$ corresponding to the number of times that $\rho$ must be applied to $\rho_{X}(w)$ before it enters $L_{1}^{\mu_{r}}\cup R_{1}^{\mu_{r}}$, that is,
@@
$\rho_{\tilde{Z}}^{J}:=\begin{cases}\theta_{\tilde{Z}}^{L_{1}^{\mu_{r}}}(\zeta)&\text{if }J\in\{L_{i}^{\gamma_{r}}\}_{i\geq 0},\vspace{0.15cm}\\ \theta_{\tilde{Z}}^{R_{1}^{\mu_{r}}}(\zeta)&\text{if }J\in\{R_{i}^{\gamma_{r}}\}_{i\geq 0}.\\ \end{cases}$
Notice that, for $J\in\{L_{i}^{\gamma_{r}}\}_{i\geq 0}$, we have $\rho_{\tilde{Z}}^{J}|_{L_{1}^{\mu_{r}}}=\theta_{\tilde{Z}}^{L_{1}^{\mu_{r}}}$, and for $J\in\{R_{i}^{\gamma_{r}}\}_{i\geq 0}$, we have $\rho_{\tilde{Z}}^{J}|_{R_{1}^{\mu_{r}}}=\theta_{\tilde{Z}}^{R_{1}^{\mu_{r}}}$. These maps are diffeomorphisms, respectively, from $L_{1}^{\mu_{r}}$ and $R_{1}^{\mu_{r}}$ onto $\gamma_{r}$, induced by $\tilde{Z}$ as defined in (8). It is important to mention that ensuring the diffeomorphic nature of these maps is the reason we chose $L_{1}^{\mu_{r}}\cup R_{1}^{\mu_{r}}$ to define $c(w)$, rather than $L_{0}^{\mu_{r}}\cup R_{0}^{\mu_{r}}$.
@@
There exists a neighborhood $U\subseteq\gamma_{r}$ of $q$ satisfying (P) for which
$0<\epsilon_{J}\leq\left\lvert{\psi^{\prime}_{J}(x)}\right\rvert\leq s<1,$
for every $J\in\cal{J}_{U}$ and $x\in\gamma_{r}$. Consequently,
$\left\lvert{\pi^{\prime}(w)}\right\rvert\geq\dfrac{1}{s}>1,$
for every $w\in U\cap\cal{W}_{r}$.
@@
Now, since $\rho_{X}$, $\rho_{\tilde{Z}}^{J}$, and $H|_{\mu_{r}}$ are diffeomorphisms on compact sets, we can set
$A:=\left(\min_{\zeta\in J}\left\lvert{(\rho_{\tilde{Z}}^{J})^{\prime}(\zeta)}\right\rvert\right)\cdot\left(\min_{z\in H(\mu_{r})}\left\lvert{(H^{-1})^{\prime}(z)}\right\rvert\right)\cdot\left(\min_{\zeta\in\mu_{r}}\left\lvert{H^{\prime}(\zeta)}\right\rvert\right)\cdot\left(\min_{x\in\gamma_{r}}\left\lvert{\rho_{X}^{\prime}(x)}\right\rvert\right)>0.$ (12)
Thus, taking the expression (11) into account, we get
@@
Before going further, it is important at this point to clarify a subtle detail. Formally, the functions of the IFS $\Psi$ are defined on the curve $\gamma_{r}$, but working with them in this context can be cumbersome. Instead, we can consider an orientation-preserving diffeomorphism $h:\gamma_{r}\to[-1,1]$ such that $h(q)=0$. By defining $\tilde{\psi}_{J}:=h\circ\psi_{J}\circ h^{-1}:[-1,1]\to h(J)$, we form the set $\tilde{\Psi}:=\big{\{}\tilde{\psi}_{J}:J\in\cal{J},J\subseteq U\big{\}}$, which is an IFS. Additionally, we have $h(L_{i}^{\gamma_{r}})\subseteq[-1,0)$ and $h(R_{i}^{\gamma_{r}})\subseteq(0,1]$. This provides a nice graphical representation of the first return map $\pi$ in Figure 4. Since Property (B3) implies that the Hausdorff dimension is invariant under bi-Lipschitz maps, which is the case for $h$, this transformation does not change any of the properties we are investigating. Therefore, from now on, we will assume, by an abuse of notation, that the functions in $\Psi$ are defined in the interval $[-1,1]$ and that the sets $J\in\cal{J}$ are subintervals of $[-1,1]$.